OpenStudy (anonymous):

Write the equation of the line, in standard form, that has an x-intercept of 2 and is parallel to 2x + y = -5.

OpenStudy (freckles):

@xo_kansasprincess_xo can you identify the slope of line given there? that what is the slope of y=-2x-5 ?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

2x-5=y? Is that the answer?

OpenStudy (freckles):

so any parallel line to the line y=-2x-5 will have the same slope

OpenStudy (freckles):

so we know we have the following so far: y=-2x+b where we need to find the y-intercept b

OpenStudy (anonymous):

2x-5=y? Is that the answer?

OpenStudy (freckles):

no y=2x-5 isn't parallel to y=-2x-5

OpenStudy (freckles):

anyways you are given the x-intercept to be 2 that is you are given a point on your line (2,0) use (2,0) to find b in the equation y=-2x+b

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Soooo now what?

OpenStudy (freckles):

I was asking you to find b in the equation y=-2x+b given (2,0)

OpenStudy (freckles):

(2,0) means when x=2 we have y=0

OpenStudy (anonymous):

0?

OpenStudy (freckles):

let me know when you have tried solving for b

OpenStudy (anonymous):

Is it 0? I just tried.

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

ok (2,0) means I'm going to replacee x with 2 and y with 0 we had y=-2x+b so now I have 0=-2(2)+b I know 2 times 2 is 4 so -2 times 2 is -4 so we have 0=-4+b

OpenStudy (freckles):

so b isn't going to be 0 because 0+(-4) isn't 0

OpenStudy (anonymous):

So how am I supposed to find out what it is?

OpenStudy (freckles):

you are suppose to find b

OpenStudy (freckles):

can you solve 0=-4+b

OpenStudy (anonymous):

HOW?

OpenStudy (anonymous):

4?

OpenStudy (freckles):

bingo you add 4 on both sides to isolate the b 0+4=-4+b+4 4=(-4+4)+b 4=0+b 4=b

OpenStudy (freckles):

so your equation that is parallel to y=-2x-5 going through (2,0) is y=-2x+4

OpenStudy (anonymous):

So now what? Lol

OpenStudy (anonymous):

It has to be in standard form

OpenStudy (freckles):

ok put y=-2x+4 in standard form then

OpenStudy (freckles):

earlier I wrote 2x+y=-5 out of standard form by subtract 2x on both sides y=-2x-5 where this is call slope-intercept form so if I wanted it back in standard form I would add 2x on both sides but anyways you have y=-2x+4 and to get this in standard form you would do what to both sides?

OpenStudy (anonymous):

2x-4=-y

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

well standard form is ax+by=c you only had one step to perform to get y=-2x+4 into that form add 2x on both sides

OpenStudy (anonymous):

I was told to multiply by -1 to make it more accurate... @hartnn told me this on a previous post of mine.

OpenStudy (anonymous):

So is the answer is -2x+4=y?

OpenStudy (freckles):

http://openstudy.com/users/xo_kansasprincess_xo#/updates/550078e9e4b04dec79279748 once you had it in -5x+3y=-44 he probably concluded to multiply both sides by -1 so he can have a positive integer c instead of a negative one but it doesn't matter unless you have a very strict definition in your algebra class for standard form we don't have extra work to do here but you sure are make it that way :p

OpenStudy (freckles):

again y=-2x+4 is slope intercept form you said you wanted the answer in standard form I suggested you add 2x on both sides you still haven't done that

OpenStudy (anonymous):

I don't understand!

OpenStudy (freckles):

y=-2x+4 I'm asking you to add 2x on both sides 2x+y=2x-2x+4 like this 2x+y=0+4 since 2x-2x=0 2x+y=4 since 0+4=4 this is actually in the form ax+by=c which is called standard form

OpenStudy (anonymous):

@sleepyjess

OpenStudy (freckles):

if you don't understand anything I said above please let me know

OpenStudy (anonymous):

I still don't understand what you're trying me to say.

OpenStudy (freckles):

Ok can you tell me which part

OpenStudy (anonymous):

I thought I had the answer. That's all I had to for the other questions.

OpenStudy (freckles):

Well I thought I gave you the answer when I added 2x on both sides for ya

OpenStudy (freckles):

but if you don't tell me what part I need to explain over it makes it kinda difficult what I need to put emphasis on

OpenStudy (freckles):

Write the equation of the line, in standard form, that has an x-intercept of 2 and is parallel to 2x + y = -5. parallel lines have the same slope x-intercept is in form (k,0) where to find the slope of 2x+y=-5 we will put this in slope intercept form to do this I will subtract 2x on both sides y=-2x-5 now this is in y=mx+b form m is the slope that is what we want to identify m=-2 the slope is -2 any parallel line will also have the slope -2 so we know linear functions are in the form y=mx+b we just found the m we want our line to have we want m to be -2 so I'm going to replace m with -2 y=-2x+b but what is b (the y-intercept) well we can find b by using the point we were given on the line we were give the x-intercept of this line is (2,0) we are given the point (2,0) is on this line (2,0) means I'm going to replace x with 2 and y with 0 in the equation y=-2x+b 0=-2(2)+b 0=-4+b since -2 times 2 is -4 4=b by adding 4 on both sides you have the equation y=-2x+4 now now to put in standard you add 2x on both sides 2x+y=2x-2x+4 2x+y = 0 +4 since 2x take away 2x is zero 2x+y = 4 since 0+4 is 4

OpenStudy (freckles):

again 2x+y=4 is in standard form because it is in the form ax+by=c where a=2 and b=1 and c=4

OpenStudy (anonymous):

That's the answer I had

OpenStudy (freckles):

it isn't the answer you showed me

OpenStudy (anonymous):

My bad lol. It's what I had written down, just didn't type it on here. Oops @freckles lol. Thanks so much:)

OpenStudy (freckles):

ok have a good day